(x^2+y^2)p +2xyq=(x+y)z


1
Expert's answer
2021-03-03T03:42:29-0500

The given equation is


     "(x\u00b2+y\u00b2) p + 2xyq = (x+y)z"


 The subsidiary equations are 


 "\dfrac{dx}{x²+y²} = \dfrac{dy}{2xy} = \dfrac{dz}{(x+y)z}"



   = "\dfrac{[zdx + zdy - (x+y)dz]}{[z(x²+y²)+2xyz- (x+y)²z]}"


   = "\dfrac{z(dx+dy) - (x+y)dz}{0}"


⇒z((dx+dy) - (x+y) )dz = 0


⇒z d(x+y) - (x+y)dz = 0


⇒(zd(x+y) - (x+y)dz) / z² = 0


⇒d((x+y) / z) = 0 


⇒(x+y) / z = c₁ 



From first two ratios we get 


"\dfrac{dx}{x²+y²} = 2\dfrac{dy}{4xy}"


 ⇒2(x²+y²) dy = 4xy dx


 ⇒2(x²+y²) dy - 4xy dx = 0


 ⇒2x² dy - 2y² dy + 4y² dy -4xy dx = 0


 ⇒2(x²-y²) dy - 2y(2xdx - 2ydy) = 0


 ⇒[(x²-y²) 2 dy - 2y d(x²-y²)]/(x²-y²)²


                 = 0/(x²-y²)


 ⇒d(2y / (x²-y²)) = 0


Integrating , we get 


  2y/(x²-y²) = c₂ 


∴ The solution is (x+y) / z = Φ(2y / (x²-y²))